On a bet, you try to remove water from a glass by
Answer #1

Use Bernoulli.
Pressure = ?1*v^2/2 = ?2*gh
?1 (air) = 1.29 kg/m^2
?2 (water) = 1000 kg/m^2
v = sqrt(2gh*?2/?1) = sqrt(2*9.81*0.016*1000/1.29) = 15.5996 m/s